Bob makes his first $1,300 deposit into an IRA earning 7.9% compounded annually on his 24th birthday and has his last $1,300 deposit on his 42nd birthday (19 equal deposits in all). With no additional deposits, the money in the IRA continues to earn 7.9% interest compounded annually until Bob retires on his 65th birthday. How much is in the IRA when Bob retires?

The amount in the IRA when Bob retires is $________

(Round to the nearest cent as needed)
Answer #1
FVAnnuity Due = c*(((1+ i)^n - 1)/i)*(1 + i )
C = Cash flow per period
i = interest rate
n = number of payments
FV= 1300*(((1+ 7.9/100)^19-1)/(7.9/100))*(1+7.9/100)
FV = 57535.67
Future value = present value*(1+ rate)^time
Future value = 57535.67*(1+0.079)^22
Future value = 306484.81